OpenStudy (anonymous):

f(x)={1-2sin(x) for x<=0} { e^-4x for x > 0} part a find f(0) part b find limit of f(x) approach 0 from left part c find limit of f (x) approach 0 from right part d id f(x) continuous at x=0

OpenStudy (anonymous):

\[f(0)\] put \(0\) in the top expresion

OpenStudy (anonymous):

since \(0\leq 0\) is true, and \(0>0\) is false, use the top one

OpenStudy (anonymous):

part b is the same as the answer to part a) since \(1-\sin(x)\) is continuous

OpenStudy (anonymous):

part c, put \(0\) in to the bottom expression

OpenStudy (anonymous):

part d if your answers to b and c are the same, then it is continuous

OpenStudy (anonymous):

part e) did you get any of that?

OpenStudy (anonymous):

f(0)={1-2sin(0) 0<=0 e-40 0>0 ?? @satellite73